A solution of Na3AsO4 is added dropwise to a solution that is 0.0602 M in Cu2+ and 0.000407 M in Ag+.

The Ksp of Cu3(AsO4)2 is 7.95e-36.
The Ksp of Ag3AsO4 is 1.03e-22.






(a) What concentration of AsO43- is necessary to begin precipitation? (Neglect volume changes.)

[AsO43-] = _________M.



(b) Which cation precipitates first?

Cu2+

Ag+     






(c) What is the concentration of AsO43- when the second cation begins to precipitate?

[AsO43-] = __________M.
Answer #1

Sol.

(a) As Ksp of Cu3(AsO4)2 is lower than the the Ksp of Ag3AsO4  

So , Cu3(AsO4)2 precipitates first and Cu2+ cation precipitates first

Reaction of Cu3(AsO4)2 :

Cu3(AsO4)2 ----> 3Cu2+ + 2AsO43-

So , Ksp = [Cu2+]3 [AsO43-]2  

7.95 × 10-36 = ( 0.0602 )3 [AsO43-]2

[AsO43-]2 = 3.64 × 10-32  

[AsO43-] = 1.91 × 10-16 M

Therefore , Concentration of AsO43- necessary to begin precipitation of Cu3(AsO4)2 is  1.91 × 10-16 M

(b)

Cu2+    cation precipitate first

(c)

As second Ag+ cation precipitate . Reaction of Ag3AsO4 :

Ag3AsO4 ----> 3Ag+ + AsO43-  

So , Ksp = [Ag+]3 [AsO43-]  

1.03 × 10-22 = ( 0.000407 )3 [AsO43-]  

[AsO43-] = 1.53 × 10-12 M

Therefore , Concentration of AsO43- when the second cation begins to precipitate is 1.53 × 10-12 M